OpenStudy (anonymous):

A uniform cylinder of radius "R" is spun about its axis to an angular velocity of w and then placed into a corner. the coefficient of friction is "k". How many turns will the cylinder make before coming to a stop?

OpenStudy (anonymous):



OpenStudy (anonymous):

just use normal equation of uniform accelertated bodies considering a uniform force of friction giving a uniform retarding acceleartion -a -a= U(co-eff of friction)*g take final W=0(omega) find the displacement....s...here will be angular displacement giving us how many radiand were swept in that time 2pi radians makes 1 turn so...find

OpenStudy (jamesj):

This is a good question. So first, let's look at how much angular momentum, L, the cylinder has initially. The moment of inertia is \[ I = mr^2 \] and hence \[ L = mr^2 \omega \] Now once the cylinder is placed in the corner there are two torques applied to it. One from the horizontal surface, one from the vertical surface. Call them tau_v and tau_h respectively. We will use the subscripts _v and _h throughout for vertical and horizontal The horizontal torque is relatively straight-forward. In magnitude it is given by \[ \tau_h = (distance \ from \ center \ of \ mass)(friction \ force \ horizontally) \] \[ = rF_{f,h} \hbox{ where $F_{f,h}$ is the horizontal friction} \] \[ = r(kN_h) \] \[ = rkmg \] Now the vertical torque in magnitude is \[ \tau_v = rF_{f,v} = rkN_v \] The challenge is to find \( N_v \). Once you have those two torques, they are equal to \[ \tau_h + \tau_v = I \alpha = I \dot{\omega} \] and you solve the equation to find when \( \omega(t) = 0 \).

OpenStudy (anonymous):

thank you both salini and JamesJ